40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 23 20:39:50 2013, Jenne, Update, LSC, PRMI sensing matrix measured at 580.1Hz SensMat_23May2013.png
    Reply  Thu May 23 22:13:12 2013, Koji, Update, LSC, PRMI sensing matrix measured at 580.1Hz 
       Reply  Fri May 24 13:41:19 2013, Jenne, Update, LSC, PRMI sensing matrix measured at 580.1Hz SensMat_23May2013_withIQ.png
          Reply  Fri May 24 14:01:59 2013, Koji, Update, LSC, PRMI sensing matrix measured at 580.1Hz 
             Reply  Fri May 24 14:40:22 2013, Jenne, Update, LSC, PRMI sensing matrix: now what? 
                Reply  Fri May 24 22:18:33 2013, Jenne, Update, LSC, PRMI sensing matrix: Got it! SensMat_23May2013_withIQ_andCorrectPhases.png
                   Reply  Tue May 28 14:52:56 2013, Jenne, Update, LSC, PRMI sensing matrix: Got it! SensMat_21May2013_withIQ_andCorrectPhases.png
                      Reply  Tue May 28 17:06:04 2013, Koji, Update, LSC, PRMI sensing matrix: Got it! 
Message ID: 8648     Entry time: Tue May 28 14:52:56 2013     In reply to: 8644     Reply to this: 8650
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRMI sensing matrix: Got it! 

Just so we have it, here is the re-analysis with the correct plot and phases for the May 21st data, taken near the violin mode:

Sensing Matrix, units = counts/meter, phase in degrees
 
            PRCL Mag   PRCL Phase    MICH Mag   MICH Phase  
AS55         9.048E+11   -22.880      2.927E+12   107.071    
REFL11       2.954E+16   -21.628      3.670E+14   123.857    
REFL33       2.757E+14  -192.608      9.186E+12    -4.037    
REFL55       2.868E+14   -82.690      1.186E+13   170.097 

SensMat_21May2013_withIQ_andCorrectPhases.png

Even though this data was taken near the violin mode (oops!), it is fairly consistent with the stuff taken a few days later at 580Hz (elog 8644). 

Neither of these is at all similar to what Kiwamu had measured a year ago (elog 6283), but we have changed many, many things since then.  He also includes an Optickle simulation, which is fairly similar to the Koji simulation in the wiki, but neither his measurements nor mine are particularly close to the simulated version.  I should think about why this is.

Also, I have fixed up the measurement scripts so that they record all of the relevant current settings / information:  Current actuator calibration, current PD demod phases, drive amplitude and drive frequency.  The "Analyze Saved Data" script has been updated to read all of this info from the files.  If you want to plot / look at any old data, open up SensMatAnalyzeSavedData in /scripts/LSC/SensingMatrix/ and put in the relevant filename that you want (which should be saved in /scripts/LSC/SensingMatrix/SensMatData/)

ELOG V3.1.3-